# Full-sky Models of Galactic Microwave Emission and Polarization at Sub-arcminute Scales for the Python Sky Model

### v1

@@ 5.2.3 Synchrotron Emission Over the Sky
In this section, we detail the validation of the new synchrotron models by comparing the power spectra with observations. For validating the PySM synchrotron total intensity models, we use the synchrotron map [^16] https://beyondplanck.science/products/files_v1 from the BeyondPlanck re-analysis of Planck LFI data (Andersen et al., 2023). This map is at a reference frequency of 30 GHz and has an angular resolution of $2^{\circ}$. We produce single-frequency total intensity maps for the different synchrotron models at 30 GHz and smoothed to a FWHM of $2^{\circ}$.
@@
We have purposely chosen an earlier release of BeyondPlanck for our analysis. Later data release versions of BeyondPlanck, or its successor CosmoGlobe, produce synchrotron intensity maps at 408 MHz (Watts et al., 2023). The PySM synchrotron models s5 and s7 are not suitable for producing synchrotron simulations at 408 MHz. This is a consequence of scaling the Haslam map from 408 MHz to 30 GHz with a constant $\beta_{s}$ for constructing the template for synchrotron intensity, and then applying a spatially variable $\beta_{s}$ to evaluate the model (see Section 3.2.1).
@@
For synchrotron polarization validation, we compare the models with Planck Revisited synchrotron polarization maps [^17] https://portal.nersc.gov/project/cmb/Planck_Revisited (Delabrouille, 2024). These are the lowest-noise full sky maps of polarized synchrotron at 30 GHz at $1^{\circ}$ resolution at the time of analysis.
@@
Figure 8: Comparison of the PySM s1 and s5 power spectra with observations over different sky fractions. The observed $TT$ power spectra are computed from the synchrotron intensity map from BeyondPlanck Release 1 (Andersen et al., 2023) at 30 GHz with a $2^{\circ}$ beam. The observed $EE$ and $BB$ power spectra are derived from the Planck Revisited (Delabrouille, 2024) polarized synchrotron map at 30 GHz with a $1^{\circ}$ beam. These $EE$ and $BB$ spectra are debiased for noise, calculated from 200 simulations, and the noise power spectra are also displayed to highlight the change of signal-to-noise ratio across the multipole range.
@@
In Figure 8 we present $TT$, $EE$ and $BB$ power spectrum comparisons for models s1 and s5. Since models s4, s6 and s7 yield nearly identical results, we display only the results for model s5. The $TT$ power spectra across all sky fractions for both models show excellent agreement with the BeyondPlanck synchrotron total intensity power spectra. For the $EE$ power spectra, model s5 demonstrates good agreement in the multipole range where the signal-to-noise is $\gtrsim 1$ for the Planck Revisited polarized synchrotron power spectra. In contrast, model s1 exhibits higher power for most multipoles compared to the Planck Revisited $EE$-synchrotron spectrum for the 80% and 60% sky fractions, although they align well for the 40% and 20% sky fractions. We also observe a distinct change in the shape of the $EE$ power spectrum for the s1 model around the injection scale, $\ell_{*}\sim 36$ (Thorne et al., 2017). The improved performance for model s5 comes from the empirical optimization of the filters in Equation 10. Both models have higher $BB$ power for 80% and 60% sky fractions.
@@
5.2.4 Dust and Synchrotron Emission in the BICEP/Keck Patch
@@
The cleanest regions of sky in the southern hemisphere are particularly crucial for ongoing and future CMB experiments aimed at measuring primordial $B$-mode polarization. The most powerful current dataset for constraining the tensor-to-scalar ratio $r$ to date comes from the BICEP/Keck (BK) experiment (Ade et al., 2021, henceforth “BK18”). This analysis includes observations up to and including the 2018 observation season, covering a $\sim 600$ square degree sky patch centered at RA 0h, Dec. $-57.5^{\circ}$, and incorporates NPIPE and WMAP data across the 23–353 GHz range from the same region. In the coming years, a collaborative effort between BICEP/Keck and the South Pole Telescope (SPT) will combine the BK maps with overlapping SPT-3G maps being used to “delens” the observed CMB $B$-modes, further tightening constraints on $r$ (The BICEP/Keck Collaboration et al., 2024). We hence provide a dedicated analysis of our models in this well-studied patch of sky.
@@
We utilize the analysis method illustrated in Figure 9, which shows PySM dust model d9 maps that have been “reobserved” to replicate the impact of the BK timestream processing and map-making pipeline. In the first panel, the d9 $Q$ and $U$ maps at 353 GHz (delta function bandpass) are masked to the BICEP3 observation region and convolved with the BICEP3 beam. The second panel displays the results after these smoothed maps are multiplied by the BICEP3 observation matrix, simulating the data reduction steps, such as filtering applied along RA scans and beam deprojection in the linear map-making process (BICEP2 Collaboration et al., 2016). This process results in output maps that appear “as observed” through the BICEP3 pipeline, with large-scale filtered out. In the third panel, the intermediate maps are multiplied by the corresponding purification matrix to remove $E$- and ambiguous-modes, leaving pure $B$-mode dominated $Q$ and $U$ maps with distinct “cross-like” and “plus-like” patterns, respectively. Finally, the inverse noise-variance apodization mask is applied.
@@
By applying the same process to other PySM models at frequencies of interest (s1, s4, s5, s6, s7 at 23, 30, 40 GHz and d1, d9, d10, d12 at 85, 150, 220, 270, 353 GHz), we generate a set of BK pipeline-propagated, pure $B$-mode model foreground maps that can be compared to BK measurements. We then calculate their $BB$ auto power spectra using the BK binning and present selected results. The color-corrected and noise-biased NPIPE $BB$ power spectrum at 353 GHz, derived from the NPIPE real and simulation maps processed using the same BK analysis method [^18] http://bicepkeck.org/bk18_2021_release.html, are also shown in this section.
@@
In Figure 10, we present the $BB$ bandpower at 353 GHz for the d1, d9, and d12 models, along with the bandpower at 23 GHz for the s1 and s5 models—these are the reference frequencies adopted in the BK analyses as well as the PySM models. The PySM power spectra are compared against the NPIPE power spectrum and BK18 foreground measurements. For dust, we show the 353 GHz dust component of the BK $BB$ spectra, derived from a reanalysis of the per-bandpower spectral decomposition at 95 GHz and 150 GHz, as shown in Figure 16 of Ade et al. (2021). For synchrotron, since the BK18 analysis does not detect a synchrotron foreground, we conduct a similar analysis at 23 GHz and instead plot the 95% upper limits on the synchrotron bandpowers. To enable a direct comparison, the PySM model and NPIPE values have been corrected for the suppression effects and partial-sky coverage which are applied by the beam, matrices and apodization mask. The figure therefore presents the full-sky $D_{\ell}$ values evaluated at delta function bandpasses.
@@
All models exhibit excess dust $BB$ power relative to the measured values in this sky patch at 353 GHz. Among them, d1 overestimates the amplitude by about a factor of three—the largest discrepancy—and it does not follow a power law in $\ell$, leading to $-\ell$ values approximately twice as high as the other models. In contrast, d9 and d12 power show smaller deviation ratios, and generally reproduce the spatial variations observed in the BK data, although d12 shows a drop in power at low $\ell$.
@@ 85 GHz 150 GHz 220 GHz 270 GHz 353 GHz
Note. — The deviation ratio of the reobserved PySM dust $BB$ spectrum to the dust component of the BK18 maximum likelihood model at each frequency.
@@
Overall, the d9 and d10 models demonstrate substantial improvements over d1, particularly in their ability to extrapolate the large-scale templates and capture the preferred small-scale power decay that the NPIPE data are unable to reveal in this specific field. However, while the discrepancies in amplitude are reduced, they still persist when compared to the BK18 measurements. Given the way in which the models are constructed, it is not to be expected that they will track reality perfectly in small clean regions. Current power spectrum-based modeling techniques require a certain degree of global averaging of power spectrum parameters, which are known to vary across the sky in small patches (Planck Collaboration IV, 2020; Córdova Rosado et al., 2024). For instance, while dust amplitude can be modulated using large-scale templates, the power-spectral tilt $\alpha_{d}$ has to be fixed to a single value for the entire sky, which is inherently unrealistic. Further refinement of the new models, which could bring this sky patch into better agreement with the BK data, will be left for future studies.
@@
Lastly, we compare the synchrotron models with the BK results, focusing on s1 and s5, as other new models yield similar outcomes. As illustrated in the lower panel of Figure 10, at 23 GHz, s5 exhibits excess power at $\ell\lesssim 50$ (or $\ell\lesssim 70$ when compared to the broad-band 95% upper limit $A_{s,\ell=80}<1.4$ $\mu\rm{K^{2}_{CMB}}$) and a robustly nonzero $\alpha_{s}$, while s1 maintains relatively constant spatial power within the 95% upper limits. This trend extends to 30 GHz and 40 GHz as well. The $-\ell$ power spectrum discrepancies between models and data are likely because the models rely on the WMAP 23 GHz map as a template only up to $\ell=38$, whereas the NPIPE 30/40 GHz maps, which favor a smaller $A_{s}$, also contribute to the BK constraints (Ade et al., 2021, Figure 21). While the BK18 data provide little constraining power on the parameters $\alpha_{s}$ and $\beta_{s}$, this comparison will be worth revisiting when new measurements from the BICEP Array telescope, especially its dedicated 30/40 GHz receiver, become available (Moncelsi et al., 2020).
5.3 Decorrelation
@@
One of the most challenging aspects of dust emission for CMB analyses is that its frequency dependence varies across the sky. If dust had the same SED everywhere, it would be sufficient to measure it at a frequency where it dominates the submillimeter emission and then subtract off all emission in lower frequency maps that is correlated with that template. The extent to which a map of dust emission at one frequency differs from a map of dust emission at a different frequency (aside from an overall normalization) is referred to as “frequency decorrelation”. Frequency decorrelation has been identified as a major uncertainty for ongoing and upcoming analyses (Ade et al., 2021).
@@
where $X$ and $Y$ can be any of $T$, $E$, or $B$ (Planck Collaboration Int. L, 2017). Here we focus on $\mathcal{R}_{\ell}^{BB}$.
The 353 and 217 GHz Planck channels have the highest sensitivity to polarized dust emission and so furnish the current best constraints on the level of dust decorrelation. Analyzing 71% of the sky over multipoles $50\leq\ell\leq 160$, Planck Collaboration XI (2020) found $\mathcal{R}_{\ell}^{BB}>0.991$ (97.5% confidence) between these two frequencies.
Figure 11: The decorrelation parameter $1-\mathcal{R}_{\ell}^{BB}$, where $\mathcal{R}_{\ell}^{BB}$ is the correlation parameter defined in Equation (16), for the d1, d10, and d12 dust models between a reference frequency of 217 GHz and variable $\nu$. $\mathcal{R}_{\ell}^{BB}$ is evaluated between $50\leq\ell\leq 160$ over the Planck 70% sky mask. The 97.5% upper limit on decorrelation derived by Planck Collaboration XI (2020) over 71% of the sky ($\mathcal{R}_{\ell}^{BB}$ > 0.991) is indicated by the black arrow at 353 GHz. The models span a large range in decorrelation level and each has a distinctive dependence on frequency.
@@
We find that all three models respect the upper limit set by Planck Collaboration XI (2020), with d12 coming closest to saturating it ($\mathcal{R}_{\ell}^{BB}(217\times 353)=0.9930$), then d10 (0.9979), then d1 (0.9993). The models thus span a range of viable levels of decorrelation. Because the spectral parameters ($T_{d}$ and $\beta_{d}$) of the d9 model do not vary across the sky, that model has no decorrelation by construction (i.e., $\mathcal{R}_{\ell}=1$).
@@
The d1 model is based on component separation with Commander that placed a Gaussian prior on $\beta_{d}$ with $\sigma_{\beta_{d}}=0.1$ (Planck Collaboration X, 2016). Therefore, most of the observed variability of the dust SED is explained in this model via fluctuations in $T_{d}$. Further, the Commander data model did not account for CIB fluctuations, and so the resulting maps of dust parameters have enhanced small-scale fluctuations from CIB contamination (see Section 5.4). In contrast, the component separation based on GNILC that led to the $T_{d}$ and $\beta_{d}$ maps used in the d10 model (Planck Collaboration Int. XLVIII, 2016) permitted large variations in $\beta_{d}$ and largely removed CIB fluctuations. Indeed, the Commander-based parameter maps have less variance in $\beta_{d}$ and more variance in $T_{d}$ than the parameter maps from the GNILC-based analysis (see Planck Collaboration Int. XLVIII, 2016, Table 1). The result is that d10 predicts much larger values of decorrelation than d1 at low frequencies where $\beta_{d}$ is the dominant driver of decorrelation, but somewhat smaller values at high frequencies where $T_{d}$ is the dominant driver.
At present, it is unclear whether d1, d9, d10, or d12 is a more accurate description of the spatial variability of $T_{d}$ and $\beta_{d}$ and thus of frequency decorrelation. Constraints on decorrelation at higher frequencies where the models diverge most sharply—such as from CCAT (CCAT-Prime Collaboration et al., 2023), Taurus (May et al., 2024), and potential balloon-borne or space-based far-infrared/submillimeter platforms like BLAST Observatory (Coppi et al., 2024), LiteBIRD (LiteBIRD Collaboration et al., 2023), PICO (Hanany et al., 2019), PIXIE (Kogut et al., 2024), and PRIMA (Dowell et al., 2024)—would enable more accurate predictions for the level of decorrelation expected at CMB frequencies. As component separation techniques continue to improve by incorporating information in both the pixel domain and the harmonic domain, and as new data sets are becoming available, it is also timely to revisit the component separation analysis with the aim of making improved $T_{d}$ and $\beta_{d}$ maps.
5.4 Extragalactic contamination

### v2

@@ V.2.3 Synchrotron Emission Over the Sky
In this section, we detail the validation of the new synchrotron models by comparing the power spectra with observations. For validating the PySM synchrotron total intensity models, we use the synchrotron map [^16] https://beyondplanck.science/products/files_v1 from the BeyondPlanck re-analysis of Planck LFI data (Andersen:2023). This map is at a reference frequency of 30 GHz and has an angular resolution of $2^{\circ}$. We produce single-frequency total intensity maps for the different synchrotron models at 30 GHz and smoothed to a FWHM of $2^{\circ}$.
@@
We have purposely chosen an earlier release of BeyondPlanck for our analysis. Later data release versions of BeyondPlanck, or its successor CosmoGlobe, produce synchrotron intensity maps at 408 MHz (Watts:2023b). The PySM synchrotron models s5 and s7 are not suitable for producing synchrotron simulations at 408 MHz. This is a consequence of scaling the Haslam map from 408 MHz to 30 GHz with a constant $\beta_{s}$ for constructing the template for synchrotron intensity, and then applying a spatially variable $\beta_{s}$ to evaluate the model (see Section III.2.1).
@@
For synchrotron polarization validation, we compare the models with Planck Revisited synchrotron polarization maps [^17] https://portal.nersc.gov/project/cmb/Planck_Revisited (Delabrouille:2024). These are the lowest-noise full sky maps of polarized synchrotron at 30 GHz at $1^{\circ}$ resolution at the time of analysis.
@@
Figure 8: Comparison of the PySM s1 and s5 power spectra with observations over different sky fractions. The observed $TT$ power spectra are computed from the synchrotron intensity map from BeyondPlanck Release 1 (Andersen:2023) at 30 GHz with a $2^{\circ}$ beam. The observed EEEE and $BB$ power spectra are derived from the Planck Revisited (Delabrouille:2024) polarized synchrotron map at 30 GHz with a $1^{\circ}$ beam. We show the $EE$ and $BB$ spectra with (colored markers) and without noise debiasing (gray markers). The noise power spectra, computed on 200 simulations, is showed by the gray shaded region.
@@
In Figure 8 we present $TT$, EEEE and $BB$ power spectrum comparisons for models s1 and s5. Since models s4, s6 and s7 yield nearly identical results, we display only the results for model s5. The $TT$ power spectra across all sky fractions for both models show excellent agreement with the BeyondPlanck synchrotron total intensity power spectra. For the $EE$ power spectra, model s5 demonstrates good agreement in the multipole range where the signal-to-noise is $\gtrsim 1$ for the Planck Revisited polarized synchrotron power spectra. In contrast, model s1 exhibits higher power for most multipoles compared to the Planck Revisited $EE$-synchrotron spectrum for the 80% and 60% sky fractions, although they align well for the 40% and 20% sky fractions. We also observe a distinct change in the shape of the $EE$ power spectrum for the s1 model around the injection scale, $\ell_{*}\sim 36$ (Thorne:2017). The improved performance for model s5 comes from the empirical optimization of the filters in Equation 10. Both models have higher $BB$ power for 80% and 60% sky fractions.
@@
V.2.4 Dust and Synchrotron Emission in the BICEP/Keck Patch
@@
The cleanest regions of sky in the southern hemisphere are particularly crucial for ongoing and future CMB experiments aimed at measuring primordial $B$-mode polarization. The most powerful current dataset for constraining the tensor-to-scalar ratio $r$ to date comes from the BICEP/Keck (BK) experiment (Ade:2021, henceforth “BK18”). This analysis includes observations up to and including the 2018 observation season, covering a $\sim 600$ square degree sky patch centered at RA 0h, Dec. $57.5^{\circ}$, and incorporates NPIPE and WMAP data across the 23–353 GHz range from the same region. In the coming years, a collaborative effort between BICEP/Keck and the South Pole Telescope (SPT) will combine the BK maps with overlapping SPT-3G maps being used to “delens” the observed CMB $B$-modes, further tightening constraints on $r$ (TheBICEP/KeckCollaboration:2024). We hence provide a dedicated analysis of our models in this well-studied patch of sky.
@@
We utilize the analysis method illustrated in Figure 9, which shows PySM dust model d9 maps that have been “reobserved” to replicate the impact of the BK timestream processing and map-making pipeline. In the first panel, the d9 $Q$ and $U$ maps at 353 GHz (delta function bandpass) are masked to the BICEP3 observation region and convolved with the BICEP3 beam. The second panel displays the results after these smoothed maps are multiplied by the BICEP3 observation matrix, simulating the data reduction steps, such as filtering applied along RA scans and beam deprojection in the linear map-making process (BICEP2Collaboration:2016). This process results in output maps that appear “as observed” through the BICEP3 pipeline, with large scales filtered out. In the third panel, the intermediate maps are multiplied by the corresponding purification matrix to remove $E$- and ambiguous-modes, leaving pure $B$-mode dominated $Q$ and $U$ maps with distinct “cross-like” and “plus-like” patterns, respectively. Finally, the inverse noise-variance apodization mask is applied.
@@
By applying the same process to other PySM models at frequencies of interest (s1, s4, s5, s6, s7 at 23, 30, 40 GHz and d1, d9, d10, d11, d12 at 85, 150, 220, 270, 353 GHz), we generate a set of BK pipeline-propagated, pure $B$-mode model foreground maps that can be compared to BK measurements. We then calculate their $BB$ auto power spectra using the BK binning and present selected results. The color-corrected and noise-biased NPIPE $BB$ power spectrum at 353 GHz, derived from the NPIPE real and simulation maps processed using the same BK analysis method [^18] http://bicepkeck.org/bk18_2021_release.html, are also shown in this section.
@@
In Figure 10, we present the $BB$ bandpower at 353 GHz for the d1, d9, and d12 models, along with the bandpower at 23 GHz for the s1 and s5 models—these are the reference frequencies adopted in the BK analyses as well as the PySM models. The PySM power spectra are compared against the NPIPE power spectrum and BK18 foreground measurements. For dust, we show the 353 GHz dust component of the BK $BB$ spectra, derived from a reanalysis of the per-bandpower spectral decomposition at 95 GHz and 150 GHz, as shown in Figure 16 of Ade:2021. For synchrotron, since the BK18 analysis does not detect a synchrotron foreground, we conduct a similar analysis at 23 GHz and instead plot the 95% upper limits on the synchrotron bandpowers. To enable a direct comparison, the PySM model and NPIPE values have been corrected for the suppression effects and partial-sky coverage which are applied by the beam, matrices and apodization mask. The figure therefore presents the full-sky $D_{\ell}$ values evaluated at delta function bandpasses.
@@
All models exhibit excess dust $BB$ power relative to the measured values in this sky patch at 353 GHz. Among them, d1 overestimates the amplitude by about a factor of three—the largest discrepancy—and it does not follow a power law in $\ell$, leading to high-$\ell$ values approximately twice as high as the other models. In contrast, d9 and d12 power show smaller deviation ratios, and generally reproduce the spatial variations observed in the BK data, although d12 shows a drop in power at low $\ell$.
@@ 85 GHz 150 GHz 220 GHz 270 GHz 353 GHz
Note. — Deviation ratios of the reobserved PySM dust $BB$ spectra to the dust component of the BK18 maximum likelihood model at each frequency. The d11 values show the mean and standard deviation derived from 100 realizations, demonstrating that d10 is indeed a representative realization from this stochastic model.
Table 6 further explores how the model amplitudes vary across frequencies, presenting the results of a single-amplitude-parameter fit between the PySM models and BK measurements down to 85 GHz. These deviation ratios are determined by fitting the nine bandpowers of the PySM models to the dust component of the BK18 maximum likelihood foreground model, which follows a modified blackbody power spectrum (with $T_{d}$ fixed at $19.6~\text{K}$) and is characterized by parameters $A_{d,\ell=80}=4.4\;\mu\rm{K^{2}_{CMB}}$, $\beta_{d}=1.5$ and $\alpha_{d}=-0.66$. d9 consistently maintains a deviation factor of around two, while d1 and d10 deviate less from the measurements at lower frequencies. This suggests that these two models, constructed using the $\beta_{d}$ maps from Commander ($\beta_{d}=1.51\pm 0.06$, planck2014-a12) and GNILC ($\beta_{d}=1.6\pm 0.1$, planck2016-XLVIII) as their respective spectral parameter templates, have too large a value of $\beta_{d}$ relative to the BK measured value of $1.49^{+0.13}_{-0.12}$ in this patch. The frequency-dependent trend observed in d12 is incompatible with a simple change of $\beta_{d}$ and likely reflects the multiple-layer behavior inherent to this model.
@@
Overall, the d9 and d10 models demonstrate substantial improvements over d1, particularly in their ability to extrapolate the large-scale templates and capture the preferred small-scale power decay that the NPIPE data are unable to reveal in this specific field. However, while the discrepancies in amplitude are reduced, notable differences still persist when compared to the BK18 measurements. Given the way in which the models are constructed, it is not to be expected that they will track reality perfectly in small clean regions, especially in the presence of the GNILC template bias we have identified. Current power spectrum-based modeling techniques require a certain degree of global averaging of power spectrum parameters, which are known to vary across the sky in small patches (planck2016-l04; CordovaRosado:2024). For instance, while dust amplitude can be modulated using large-scale templates, the power-spectral tilt $\alpha_{d}$ has to be fixed to a single value for the entire sky, which is inherently unrealistic. Further refinement of the new models, which could bring this sky patch into better agreement with the BK data, will be left for future studies.
@@
Lastly, we compare the synchrotron models with the BK results, focusing on s1 and s5, as other new models yield similar outcomes. As illustrated in the lower panel of Figure 10, at 23 GHz, s5 exhibits excess power at $\ell\lesssim 50$ (or $\ell\lesssim 70$ when compared to the broad-band 95% upper limit $A_{s,\ell=80}<1.4$ $\mu\rm{K^{2}_{CMB}}$) and a robustly nonzero $\alpha_{s}$, while s1 maintains relatively constant spatial power within the 95% upper limits. This trend extends to 30 GHz and 40 GHz as well. The low-$\ell$ power spectrum discrepancies between models and data are likely because the models rely on the WMAP 23 GHz map as a template only up to $\ell=38$, whereas the NPIPE 30/40 GHz maps, which favor a smaller $A_{s}$, also contribute to the BK constraints (Ade:2021, Figure 21). While the BK18 data provide little constraining power on the parameters $\alpha_{s}$ and $\beta_{s}$, this comparison will be worth revisiting when new measurements from the BICEP Array telescope, especially its dedicated 30/40 GHz receiver, become available (Moncelsi:2020).
V.3 Decorrelation
@@
One of the most challenging aspects of dust emission for CMB analyses is that its frequency dependence varies across the sky. If dust had the same SED everywhere, it would be sufficient to measure it at a frequency where it dominates the submillimeter emission and then subtract off all emission in lower frequency maps that is correlated with that template. The extent to which a map of dust emission at one frequency differs from a map of dust emission at a different frequency (aside from an overall normalization) is referred to as “frequency decorrelation”. Frequency decorrelation has been identified as a major uncertainty for ongoing and upcoming analyses (Ade:2021).
@@
where $X$ and $Y$ can be any of $T$, $E$, or $B$ (planck2016-L). Here we focus on $\mathcal{R}_{\ell}^{BB}$.
The 353 and 217 GHz Planck channels have the highest sensitivity to polarized dust emission and so furnish the current best constraints on the level of dust decorrelation. Analyzing 71% of the sky over multipoles $50\leq\ell\leq 160$, planck2016-l11A found $\mathcal{R}_{\ell}^{BB}>0.991$ (97.5% confidence) between these two frequencies.
Figure 11: The decorrelation parameter $1-\mathcal{R}_{\ell}^{BB}$, where $\mathcal{R}_{\ell}^{BB}$ is the correlation parameter defined in Equation (16), for the d1, d10, and d12 dust models between a reference frequency of 217 GHz and variable $\nu$. $\mathcal{R}_{\ell}^{BB}$ is evaluated between $50\leq\ell\leq 160$ over the Planck 70% sky mask. The 97.5% upper limit on decorrelation derived by planck2016-l11A over 71% of the sky ($\mathcal{R}_{\ell}^{BB}$ > 0.991) is indicated by the black arrow at 353 GHz. The models span a large range in decorrelation level and each has a distinctive dependence on frequency.
@@
We find that all three models respect the upper limit set by planck2016-l11A, with d12 coming closest to saturating it ($\mathcal{R}_{\ell}^{BB}(217\times 353)=0.9930$), then d10 (0.9979), then d1 (0.9993). The models thus span a range of viable levels of decorrelation. Because the spectral parameters ($T_{d}$ and $\beta_{d}$) of the d9 model do not vary across the sky, that model has no decorrelation by construction (i.e., $\mathcal{R}_{\ell}=1$).
@@
The d1 model is based on component separation with Commander that placed a Gaussian prior on $\beta_{d}$ with $\sigma_{\beta_{d}}=0.1$ (planck2014-a12). Therefore, most of the observed variability of the dust SED is explained in this model via fluctuations in $T_{d}$. Further, the Commander data model did not account for CIB fluctuations, and so the resulting maps of dust parameters have enhanced small-scale fluctuations from CIB contamination (see Section V.4). In contrast, the component separation based on GNILC that led to the $T_{d}$ and $\beta_{d}$ maps used in the d10 model (planck2016-XLVIII) permitted large variations in $\beta_{d}$ and largely removed CIB fluctuations. Indeed, the Commander-based parameter maps have less variance in $\beta_{d}$ and more variance in $T_{d}$ than the parameter maps from the GNILC-based analysis (see planck2016-XLVIII, Table 1). The result is that d10 predicts much larger values of decorrelation than d1 at low frequencies where $\beta_{d}$ is the dominant driver of decorrelation, but somewhat smaller values at high frequencies where $T_{d}$ is the dominant driver.
At present, it is unclear whether d1, d9, d10, or d12 is a more accurate description of the spatial variability of $T_{d}$ and $\beta_{d}$ and thus of frequency decorrelation. Constraints on decorrelation at higher frequencies where the models diverge most sharply—such as from CCAT (CCAT-PrimeCollaboration:2023), SPIDER-2 (Shaw:2024), Taurus (May:2024), and potential balloon-borne or space-based far-infrared/submillimeter platforms like BLAST Observatory (Coppi:2024), LiteBIRD (LiteBIRDCollaboration:2023), PICO (Hanany:2019), PIXIE (Kogut:2025), and PRIMA (Dowell:2024)—would enable more accurate predictions for the level of decorrelation expected at CMB frequencies. As component separation techniques continue to improve by incorporating information in both the pixel domain and the harmonic domain, and as new data sets are becoming available, it is also timely to revisit the component separation analysis with the aim of making improved $T_{d}$ and $\beta_{d}$ maps.
V.4 Extragalactic contamination